OpenStudy (anonymous):

a sweet factory makes 25000 lollipops every year. 12% of them are blackcurrant and 43% of them are orange flavoured. The rest of them are strawberry flavoured... how many strawberry flavoured are made throughout the year?? Is the answer 11250??

jimthompson5910 (jim_thompson5910):

getting the same answer 12+43 = 55% are blackcurrant or orange so 100 - 55 = 45% are strawberry 0.45*25,000 = 11,250

OpenStudy (anonymous):

yeyeeee, thankyou... i wasnt to sure so just thought i'd check with you brainboxes ;)

jimthompson5910 (jim_thompson5910):

you're welcome and congrats on getting the right answer

OpenStudy (anonymous):

hehehe, it wasnt to hard but it took a bit of thinking :P

jimthompson5910 (jim_thompson5910):

glad you found it easy, it's not too bad once you get used to percentages and such

OpenStudy (anonymous):

:)